OpenStudy (anonymous):

solve the systems of equations by graphing then classify the system 6x-y=25 6x+7y=-31

OpenStudy (anonymous):

Y=-7, X=3

OpenStudy (anonymous):

Am i clear to you?

OpenStudy (anonymous):

6x-y=25 -(6x+7y=-3) - 8y=56 y=-7 x=3